OpenStudy (anonymous):

what is the integral of (e^(1/x^2)) / (x^3) dx?

OpenStudy (anonymous):

\[\int\limits \frac{e^{\frac{1}{x^2}}}{x^3} \, dx=-\frac{1}{2} e^{\frac{1}{x^2}}+c \]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Yes! Finally